# Continuous Ensemble Weather Forecasting with Diffusion models

### v1

@@ 1 Introduction
Forecasting of physical systems over both space and time is a crucial problem with plenty of real-world applications, including in the earth sciences, transportation, and energy systems. A prime example of this is weather forecasting, which billions of people depend on daily to plan their activities. Weather forecasting is also crucial for making informed decisions in areas such as agriculture, renewable energy production, and safeguarding communities against extreme weather events. Current Numerical weather prediction (NWP) systems predict the weather using complex physical models and large supercomputers (Bauer et al., 2015). Recently Machine learning weather prediction (MLWP) models have emerged, rivaling the performance of existing NWP systems. These models are not physics-based but data-driven and have been made possible thanks to advancements in deep learning. By analyzing patterns from vast amounts of meteorological data (Hersbach et al., 2020), MLWP models now predict the weather with the same accuracy as global operational NWP models in a fraction of the time (Kurth et al., 2023; Lam et al., 2023; Bi et al., 2023).
@@
Following the success of deterministic MLWP models, the focus of the field has increasingly shifted towards probabilistic modeling. The probabilistic models generate samples of possible future weather trajectories. By drawing many such samples, referred to as ensemble members, it is possible to generate an ensemble forecast for quantifying forecast uncertainty and detecting extreme events (Leutbecher & Palmer, 2008). Sampling forecasts from deep generative models also address the blurriness often observed in predictions from deterministic MLWP, yielding forecasts that better preserve the variability of the modeled entities. A popular class of deep generative models used for probabilistic MLWP are diffusion models (Ho et al., 2020; Price et al., 2024; Lang et al., 2024; Shi et al., 2024). While these models generate accurate and realistic looking forecasts, they are computationally expensive due to requiring multiple sequential forward passes through the neural network to generate a sample. Moreover, they are often applied iteratively to roll out longer forecasts (Price et al., 2024), which exacerbates the computational issue. Naively switching out this iterative rollout to directly forecasting each future time step does not result in trajectories that are consistent over time. The auto-regressive rollout additionally puts some limitations on the temporal resolution of the forecast. Taking too small timesteps results in large accumulation of error over time (Bi et al., 2023), which forces existing models to resort to a temporal resolution of 12 h Price et al. (2024); Lang et al. (2024). However, in many situations it is crucial to obtain probabilistic forecasts at a much higher temporal resolution. This is true not least when the forecasts are used as decision support in extreme weather situations and for capturing rapidly changing weather events.
@@
We propose a continuous forecasting diffusion model that takes lead time as input and forecasts the future weather state in a single step, while maintaining a temporally consistent trajectory for each ensemble member. [^1] Our code will be made openly available at publication. https://github.com/martinandrae/Continuous-Ensemble-Forecasting This enables both autoregressive and direct forecasting within a single model, improves the accuracy compared to purely autoregressive models at high temporal resolutions, and enables forecasting at arbitrary (non-equidistant) lead times throughout the forecast trajectory. To generate ensemble forecasts, the model uses a lead-time-dependent deterministic ODE-solver to solve the probability flow ODE starting in different pure noise samples. To ensure temporal consistency, we correlate the driving noises for the different lead times, e.g., by using a single noise sample for all timesteps, as illustrated in fig. 1, which enables generating a continuous trajectory for each member. This enables parallel sampling of individual ensemble members without the use of any multi-step loss functions, significantly enhancing the efficiency of ensemble forecasting.
TimeRefer to captionInitial conditionRefer to caption24 hoursRefer to caption6 hoursRefer to caption1 hourRefer to captionFixed noiseDeterministic ODE-solver
Figure 1: Diagram of Continuous Ensemble Forecasting.
@@ 2 Related Work
Unlike in NWP, where stability conditions dictate the time step, MLWP models are free to predict at any temporal resolution. Still, the most common approach is to learn to forecast a single short time-step (6h) and iterate this process until the desired lead time (Lam et al., 2023; Chen et al., 2023b). Although intuitive, this process can lead to error accumulation and is impossible to parallelize due to its sequential nature (Bi et al., 2023). Multi-step losses have been shown to reduce the error accumulation for deterministic (Lam et al., 2023) and latent-variable based models (Oskarsson et al., 2024), but are not trivially implemented in diffusion models. Taking longer timesteps (24h) has been shown to give better results (Couairon et al., 2024; Bi et al., 2023), but comes at the loss of temporal resolution. Bi et al. (2023) resolves this by training multiple models to forecast different lead times, which are then combined in different ways to reach the lead times of interest. Nguyen et al. (2023a; b) use a similar setup, but train a single model taking the forecast lead time as an input. This continuous forecasting parallelizes the prediction of the fine temporal scales, but has so far only been applied to deterministic models. Other approaches have tried to learn fully time-continuous dynamics by using an ODE to generate forecasts (Verma et al., 2024; Saleem et al., 2024; Rühling Cachay et al., 2023; Kochkov et al., 2024).
@@
Outside of MLWP, diffusion models have also been applied to forecasting other spatio-temporal processes (Yang et al., 2024). Notable examples include turbulent flow simulation (Kohl et al., 2024; Rühling Cachay et al., 2023) and PDE solving (Lippe et al., 2023). Yang & Sommer (2023) apply diffusion models conditioned on the prediction lead time to a specific floating-smoke fluid field, but do not consider ensemble forecasting. Another alternative to autoregressive rollouts is the DYfussion framework (Rühling Cachay et al., 2023), where stochastic interpolation and deterministic forecasting is combined into a diffusion-like model. The method allows for forecasting at arbitrary temporal resolution, but still requires sequential computations for sampling the prediction. DYffusion has been successfully applied to climate modeling (Cachay et al., 2024), but not weather forecasting.
@@ 3 Background
Similarly to Price et al. (2024), we model $p(X[k{+}1]|X[k{-}M{:}k])$ using a conditional diffusion model (Ho et al., 2020; Song et al., 2021; Karras et al., 2022). Diffusion models generate samples by iteratively transforming noise into data. To forecast a future weather state $X[k{+}1]$ given $X[k{-}M{:}k]$ we sample a latent noise variable $Z\sim p_{\text{noise}}$ and iteratively transform it until it resembles a sample from $p(X[k{+}1]|X[k{-}M{:}k])$. We consider the SDE formulation of diffusion models presented by Karras et al. (2022), which separates training and sampling. Sampling can then be done by solving the probability flow ODE
$\text{d}\bm{z}=-\sigma S_{\theta}(\bm{z};X[k{-}M{:}k],\sigma)\text{d}\sigma$ (2)
starting in $Z=\bm{z}(0)\sim p_{\text{noise}}$ and ending in $X[k{+}1]=\bm{z}(1)$. Here $S_{\theta}$ is the neural network trained to match the score function $S$ through the denoising training objective as presented by Karras et al. (2022). The noise level $\sigma$ is fed to $S_{\theta}$ in each layer as a Fourier embedding similar to Karras et al. (2022). Repeating this process for different noise samples $\bm{z}(0)$ gives an ensemble of forecasts. For a more detailed description of diffusion models see appendix B.
@@ 4 Continuous Ensemble Forecasting
Iterative forecasting is simple to train but can suffer from error accumulation at long horizons. Consider again the general problem of sampling from $p(X(\mathcal{T})|X(\Omega))$. In continuous forecasting, a single-step prediction is given by conditioning on the lead time $t\in\mathcal{T}$ and training a conditional score network $S_{\theta}(\bm{z};X(\Omega),t,\sigma)$ to simulate directly from the marginal distribution $p(X({t})|X(\Omega),t)$. This does not require setting a fixed $\delta$, making the setup more flexible. The lead time $t$ is added to the conditioning arguments for clarity and can be passed to the network in the same way as the noise level $\sigma$. While this allows sampling a distribution of states $X(t)$ at each time $t\in\mathcal{T}$, combining these naively does not result in a trajectory $X(\mathcal{T})$. This is because $X(t)$ are samples from the marginal distributions $p(X(t)|X(\Omega),t)$ and not the joint trajectory distribution $p(X(\mathcal{T})|X(\Omega))$. We propose Continuous Ensemble Forecasting, a novel method of combining samples from $p(X(t)|X(\Omega),t)$ into forecast trajectories $X(\mathcal{T})$ without resorting to autoregressive predictions.
The core idea in our method is to control the source of randomness. To sample from $p(X(t)|X(\Omega),t)$, we sample some noise $Z\sim p_{\text{noise}}$ and feed it to an ODE-solver that solves the probability flow ODE, giving us a forecast $X(t)$ for lead time $t$. Since the ODE-solver is deterministic, the randomness is limited to the noise initialization $Z$. If we freeze the noise, the ODE-solver becomes a deterministic map $f_{\theta}^{Z}:\mathcal{X}^{|\Omega|}\times\mathcal{T}\to\mathcal{X}$ parameterized by the neural network $S_{\theta}$. Applying this map to previous states $X(\Omega)$ and a time $t\in\mathcal{T}$ gives a forecast $X(t)$. Repeating this for several $t_{1},\dots,t_{N}\subset\mathcal{T}$ gives a sequence of forecasts $X(\{t_{i}\}_{i=1}^{N})$. Extending this to all $t\in\mathcal{T}$, we can construct a trajectory $X(\mathcal{T})=f_{\theta}^{Z}(X(\Omega),\mathcal{T})$. We treat this as a sample from $p(X(\mathcal{T})|X(\Omega))$ and propose to use Algorithm 1 to sample such trajectories.
@@
2:sample $\{z^{k}\}_{k=1}^{n_{\text{ens}}}\sim\mathcal{N}(0,\mathbf{I})$
3:for all $i\in\{1,\dots,N\},k\in\{1,\dots,n_{\text{ens}}\}$ do $\triangleright$ Can be done fully in parallel for all $k$ and $i$
4:     $x^{k}_{i}\leftarrow\textsc{ode-solver}(z^{k},t_{i};x(\Omega),S_{\theta})$ return $\{x^{k}_{i}\}^{k=1:n_{\text{ens}}}_{i=1:N}$
@@ 4.1 Mathematical motivation
The key insight in motivating our method is the identification of the latent noise space $(\mathcal{X},p_{\text{noise}})$ with the solution space $(\mathcal{F},\mu)$. In our method, the sampling algorithm can be represented by a parameterized function $f_{\theta}^{Z}$, where we have frozen the noise $Z\sim p_{\text{noise}}$. Under some regularity conditions, this function is uniquely defined by $\theta$ and the ODE-solver for any given $Z$. Thus, as illustrated in fig. 2, our setting mirrors the theoretical setting. Given sufficient data and model capacity, the neural network $S_{\theta}$ matches the score function $S$. Consequently, the distribution of the functions $f_{\theta}^{Z}$ should mirror that of the solutions $f^{i}$. Since $f^{i}$ describes a possible evolution of weather, it has to be continuous as a function of time. To ensure that the generated trajectories are also continuous in $t$, regularity conditions need to be imposed on $S_{\theta}$ to ensure that it depends smoothly on the lead time. While we do not impose any restrictions on the network, we show empirically in sec. 5.1 that this is satisfied in practice, and leave a formal analysis for future work.
@@ 4.2 Autocorrelated noise
As mentioned above, under regularity conditions, we expect the solution map $f_{\theta}^{Z}(X(\Omega),t)$ to be continuous in $t$. A shortcoming of this approach, however, is that it also constrains the trajectories to be conditionally deterministic, conditionally on the solution $f_{\theta}^{Z}(X(\Omega),t^{\star})$ at any fixed time point $t^{\star}\in\mathcal{T}$. Specifically, consider the conditional distribution $p(X(t)\mid X(t^{\star}),X(\Omega),t)$ for $t,t^{\star}\in\mathcal{T}$ under the proposed model. Conditionally on $X(t^{\star})=f_{\theta}^{Z}(X(\Omega),t^{\star})$ we can, conceptually, invert the ODE which generated $X(t^{\star})$ to recover the driving noise $Z$. Now, if $X(t)=f_{\theta}^{Z}(X(\Omega),t)$ is generated using the exact same noise variable, we find that $p(X(t)\mid X(t^{\star}),X(\Omega),t)$ is a Dirac point mass concentrated at $f_{\theta}^{Z}(X(\Omega),t)$. This violates the assumed stochasticity of the dynamical process that we are modelling.
To address this shortcoming, a simple extension of the proposed model is to replace the fixed $Z$ with a stochastic process $Z(\mathcal{T})$ with continuous sample trajectories. The stochastic process is chosen to be stationary with marginal distribution $Z(t)\sim p_{\text{noise}},\forall t\in\mathcal{T}$. This ensures that the generative model at any fixed $t$ is probabilistically equivalent with the fixed noise setting. Specifically, no changes to the training algorithm are needed since the training is based solely on time marginals. Allowing temporal stochasticity in the driving noise process results in a non-deterministic relationship between the states at different lead times, while we can keep the temporal consistency by ensuring that the driving process is sufficiently autocorrelated. A simple choice is to select $Z(\mathcal{T})$ as an Ornstein-Uhlenbeck process (Gillespie, 1996), $dZ(t)=-\rho Z(t)dt+\sqrt{2\rho}dW(t),$ for some correlation parameter $\rho>0$ and $W(t)$ being Brownian motion. This is what we use in Algorithm 2. Alternatively, we can use a Gaussian process with stronger smoothness properties, e.g., with a squared exponential kernel. This could prove useful for formally proving time continuity of the resulting trajectories, but we leave such an analysis for future work.
@@
2:sample $\{z^{k}_{1}\}_{k=1}^{n_{\text{ens}}}\sim\mathcal{N}(0,\mathbf{I})$
@@
4:     sample $\{\nu^{k}_{i}\}_{k=1}^{n_{\text{ens}}}\sim\mathcal{N}(0,\mathbf{I})$
5:     $z_{i}^{k}\leftarrow\exp(-\rho(t_{i}-t_{i-1}))z^{k}_{i-1}+\sqrt{1-\exp(-2\rho(t_{i}-t_{i-1}))}\nu^{k}_{i}$
6:for all $i\in\{1,\dots,N\},k\in\{1,\dots,n_{\text{ens}}\}$ do $\triangleright$ Can be done fully in parallel for all $k$ and $i$
7:     $x^{k}_{i}\leftarrow\textsc{ode-solver}(z^{k}_{i},t_{i};x(\Omega),S_{\theta})$ return $\{x^{k}_{i}\}^{k=1:n_{\text{ens}}}_{i=1:N}$

### v2

@@ 1 Introduction
Forecasting of physical systems over both space and time is a crucial problem with plenty of real-world applications, including in earth sciences, transportation, and energy systems. A prime example of this is weather forecasting, which billions of people depend on daily to plan their activities. Weather forecasting is also crucial for making informed decisions in areas such as agriculture, renewable energy production, and safeguarding communities against extreme weather events. Current Numerical weather prediction (NWP) systems predict the weather using complex physical models and large supercomputers (Bauer et al., 2015). Recently Machine learning weather prediction (MLWP) models have emerged, rivaling the performance of existing NWP systems. These models are not physics-based but data-driven and have been made possible thanks to advancements in deep learning. By analyzing patterns from vast amounts of meteorological data (Hersbach et al., 2020), MLWP models now predict the weather with the same accuracy as global operational NWP models in a fraction of the time (Kurth et al., 2023; Lam et al., 2023; Bi et al., 2023).
@@
Following the success of deterministic MLWP models, the focus of the field has increasingly shifted towards probabilistic modeling. The probabilistic models generate samples of possible future weather trajectories. By drawing many such samples, referred to as ensemble members, it is possible to generate a set of possible forecasts, referred to as an ensemble forecast, for quantifying forecast uncertainty and detecting extreme events (Leutbecher & Palmer, 2008). Sampling forecasts from deep generative models also address the blurriness often observed in predictions from deterministic MLWP, yielding forecasts that better preserve the variability of the modeled entities. A popular class of deep generative models used for probabilistic MLWP are diffusion models (Price et al., 2025; Lang et al., 2024; Larsson et al., 2025). While these models generate accurate and realistic looking forecasts, they are computationally expensive due to requiring multiple sequential forward passes through the neural network to generate a sample. Moreover, they are often applied iteratively to roll out longer forecasts (Price et al., 2025), which exacerbates the computational issue. Naively switching out this iterative rollout to directly forecasting each future timestep does not result in trajectories that are consistent over time. The auto-regressive rollout additionally puts some limitations on the temporal resolution of the forecast. Taking too small timesteps results in large accumulation of error over time (Bi et al., 2023), which forces existing models to resort to a temporal resolution of 12h Price et al. (2025). However, in many situations it is crucial to obtain probabilistic forecasts at a much higher temporal resolution. This is true not least when the forecasts are used as decision support in extreme weather situations and for capturing rapidly changing weather events.
@@
We propose a continuous forecasting diffusion model that takes lead time as input and forecasts the future weather state in a single step, while maintaining a temporally consistent trajectory for each ensemble member.  [^1] The code is available at https://github.com/martinandrae/Continuous-Ensemble-Forecasting This enables both autoregressive and direct forecasting within a single model, improves the accuracy compared to purely autoregressive models at high temporal resolutions, and enables forecasting at arbitrary (non-equidistant) lead times throughout the forecast trajectory. To generate ensemble forecasts, the model solves the lead-time-dependent probability flow ODE starting in different pure noise samples. To ensure temporal consistency, we correlate the driving noises for the different lead times, e.g., by using a single noise sample for all timesteps, as illustrated in fig. 1, which enables generating a continuous trajectory for each member. This enables parallel sampling of individual ensemble members, bypassing the need for multi-step loss functions, and thus accelerating ensemble forecasting.
[…]
Figure 1: Our proposed framework, Continuous Ensemble Forecasting, generates ensemble weather forecasts using a conditional diffusion model. The model takes lead time as input and forecasts the future weather state in a single step, e.g. the forecast at 24 hours is generated directly from the initial condition, without seeing the intermediate predictions. To ensure temporal consistency, we correlate the driving noises for the different lead times. This can be done by fixing the noise, or by defining a stochastic noise process. Repeating the procedure for different starting noise gives an ensemble of forecasts. Using this framework we can generate 10 day forecasts with 1 hour resolution without sacrificing performance.
@@ 2 Related Work
Unlike in NWP, where stability conditions dictate the timestep, MLWP models are free to predict at any temporal resolution. Still, the most common approach is to learn to forecast a single short timestep (6h) and iterate this process until the desired lead time (Lam et al., 2023; Chen et al., 2023b). Although intuitive, this process can lead to error accumulation and is impossible to parallelize due to its sequential nature (Bi et al., 2023). Multi-step losses have been shown to reduce the error accumulation for deterministic (Lam et al., 2023) and latent-variable based models (Oskarsson et al., 2024), but are not trivially implemented in diffusion models. Taking longer timesteps (24h) has been shown to give better results (Couairon et al., 2024; Bi et al., 2023), but comes at the loss of temporal resolution. Bi et al. (2023) resolves this by training multiple models to forecast different lead times, which are then combined in different ways to reach the lead times of interest. Nguyen et al. (2023; 2025) use a similar setup, but train a single model taking the forecast lead time as an input. This approach, called continuous forecasting, parallelizes the prediction of the fine temporal scales, but has so far only been applied to deterministic models. Other approaches have tried to learn fully time-continuous dynamics by using an ODE to generate forecasts (Verma et al., 2024; Saleem et al., 2024; Rühling Cachay et al., 2023; Kochkov et al., 2024).
@@
Outside of MLWP, diffusion models have also been applied to forecasting other spatio-temporal processes (Yang et al., 2024). Notable examples include turbulent flow simulation (Kohl et al., 2024; Rühling Cachay et al., 2023) and PDE solving (Lippe et al., 2023). Yang & Sommer (2023) apply diffusion models conditioned on the prediction lead time to a specific floating-smoke fluid field, but do not consider ensemble forecasting. Another alternative to autoregressive rollouts is the DYffusion framework (Rühling Cachay et al., 2023), where stochastic interpolation and deterministic forecasting is combined into a diffusion-like model. The method allows for forecasting at arbitrary temporal resolution, but still requires sequential computations for sampling the prediction. To get a probabilistic model, they introduce a layer dropout term in the interpolator that they keep on during inference. This makes the performance sensitive to the dropout rate, which can not be changed without retraining both the interpolator and forecasting networks. Further, the interpolation gives no guarantee of temporal continuity of trajectories, and since its trained with a mean squared error loss, is prone to blurring similar to deterministic forecasting models. DYffusion has been successfully applied to climate modeling (Rühling Cachay et al., 2025), but not weather forecasting.
@@ 3 Background
Similarly to Price et al. (2025), we model $p(X[k{+}1]|X[k{-}M{:}k])$ using a conditional diffusion model (Ho et al., 2020; Song et al., 2021; Karras et al., 2022). Diffusion models generate samples by iteratively transforming noise into data. To forecast a future weather state $X[k{+}1]$ given $X[k{-}M{:}k]$ we sample random noise from $p_{\text{noise}}$ and iteratively transform it until it resembles a sample from $p(X[k{+}1]|X[k{-}M{:}k])$. We consider the SDE formulation of diffusion models presented by Karras et al. (2022), but remark that our framework generalizes to any diffusion or flow matching framework based on stochastic differential equations. Sampling can then be done by solving the probability flow ODE
${z}(0)={z}(1)-\int_{0}^{1}\dot{\sigma}_{s}\sigma_{s}S_{\theta}\left({z}(s),\sigma_{s};X[k{-}M{:}k]\right)\mathrm{d}s,\quad{z}(1)\sim p_{\text{noise}}$ (2)
starting in pure noise and ending in our forecast ${z}(0)\sim p(X[k{+}1]|X[k{-}M{:}k])$. The ODE-solver can be chosen freely, we employ the second-order Heun’s method. Here $S_{\theta}$ is the neural network trained to match the score function $\nabla_{z}\log p_{s}$ through the denoising training objective as presented by Karras et al. (2022). Similar to (Karras et al., 2022), we choose $\sigma_{s}=s$ and feed the noise level $\sigma$ to $S_{\theta}$ in each layer as a Fourier embedding. Repeating this process for different noise samples ${z}(1)$ gives an ensemble of forecasts. For a more detailed description of diffusion models see appendix B.
@@ 4 Continuous Ensemble Forecasting
Autoregressive forecasting models are simple to train but can suffer from error accumulation at long horizons. Consider again the general problem of sampling from $p(X(\mathcal{T})|X(\Omega))$. In continuous forecasting, a single-step prediction is given by conditioning on the lead time $t\in\mathcal{T}$ and training a conditional score network $S_{\theta}({z},\sigma;X(\Omega),t)$ to simulate directly from the marginal distribution $p(X({t})|X(\Omega),t)$. This does not require setting a fixed $\delta$, making the setup more flexible. The lead time $t$ is added to the conditioning arguments for clarity and can be passed to the network in the same way as the noise level $\sigma$. While this allows sampling a distribution of states $X(t)$ at each time $t\in\mathcal{T}$, combining these naively does not result in a trajectory $X(\mathcal{T})$. This is because $X(t)$ are samples from the marginal distributions $p(X(t)|X(\Omega),t)$ and not the joint trajectory distribution $p(X(\mathcal{T})|X(\Omega))$. We propose Continuous Ensemble Forecasting, a novel method of combining samples from $p(X(t)|X(\Omega),t)$ into forecast trajectories $X(\mathcal{T})$ without resorting to autoregressive predictions.
The core idea in our method is to control the source of randomness. To sample from $p(X(t)|X(\Omega),t)$, we sample some noise $Z\sim p_{\text{noise}}$ and feed it to an ODE-solver that solves the probability flow ODE in eq. 2, giving us a forecast $X(t)$ for lead time $t$. Since the ODE-solver is deterministic, the randomness is limited to the noise initialization $Z$. If we freeze the noise, the ODE-solver becomes a deterministic map $f_{\theta}^{Z}:\mathcal{X}^{|\Omega|}\times\mathcal{T}\to\mathcal{X}$ parameterized by the neural network $S_{\theta}$. Applying this map to previous states $X(\Omega)$ and a time $t\in\mathcal{T}$ gives a forecast $X(t)$. Repeating this for several $t_{1},\dots,t_{N}\subset\mathcal{T}$ gives a sequence of forecasts $X(\{t_{i}\}_{i=1}^{N})$. Extending this to all $t\in\mathcal{T}$, we can construct a trajectory $X(\mathcal{T})=f_{\theta}^{Z}(X(\Omega),\mathcal{T})$. We treat this as a sample from $p(X(\mathcal{T})|X(\Omega))$ and propose to use Algorithm 1 to sample such trajectories.
@@
2:sample $\{z^{j}\}_{j=1}^{n_{\text{ens}}}\sim\mathcal{N}(0,\mathbf{I})$
3:for all $i\in\{1,\dots,N\},j\in\{1,\dots,n_{\text{ens}}\}$ do $\triangleright$ Can be done fully in parallel for all $j$ and $i$
4:     $x^{j}_{i}\leftarrow\textsc{{Probability-Flow-Solver}}(z^{j},t_{i};x(\Omega),S_{\theta})$ return $\{x^{j}_{i}\}^{j=1:n_{\text{ens}}}_{i=1:N}$
@@ 4.1 Mathematical motivation
The key insight in motivating our method is the identification of the latent noise space $(\mathcal{X},p_{\text{noise}})$ with the solution space $(\mathcal{F},\mu)$. In our method, the sampling algorithm can be represented by a parameterized function $f_{\theta}^{Z}$, where we have frozen the noise $Z\sim p_{\text{noise}}$. Under the regularity conditions specified in appendix E, this function is uniquely defined by $\theta$ and the ODE-solver for any given $Z$. Thus, as illustrated in fig. 2, our setting mirrors the theoretical setting. Given sufficient data and model capacity, the neural network $S_{\theta}$ matches the score function $S$. Consequently, the distribution of the functions $f_{\theta}^{Z}$ should mirror that of the solutions $f^{i}$. Since $f^{i}$ describes a possible evolution of weather, it has to be continuous as a function of time. To ensure that the generated trajectories are also continuous in $t$, regularity conditions need to be imposed on $S_{\theta}$ to ensure that it depends smoothly on the lead time. In appendix E, we provide sufficient conditions and a proof of this property. We also show empirically in sec. 5.1 that this is satisfied in practice.
@@ 4.2 Autocorrelated noise
As mentioned above, under regularity conditions on $S_{\theta}$, we expect the solution map $f_{\theta}^{Z}(X(\Omega),t)$ to be continuous in $t$. A shortcoming of this approach, however, is that it also constrains the trajectories to be conditionally deterministic, conditionally on the solution $f_{\theta}^{Z}(X(\Omega),t^{\star})$ at any fixed time point $t^{\star}\in\mathcal{T}$. Specifically, consider the conditional distribution $p(X(t)\mid X(t^{\star}),X(\Omega),t)$ for $t,t^{\star}\in\mathcal{T}$ under the proposed model. Conditionally on $X(t^{\star})=f_{\theta}^{Z}(X(\Omega),t^{\star})$ we can, conceptually, invert the ODE which generated $X(t^{\star})$ to recover the driving noise $Z$. Now, if $X(t)=f_{\theta}^{Z}(X(\Omega),t)$ is generated using the exact same noise variable, we find that $p(X(t)\mid X(t^{\star}),X(\Omega),t)$ is a Dirac point mass concentrated at $f_{\theta}^{Z}(X(\Omega),t)$. This violates the assumed stochasticity of the dynamical process that we are modelling.
To address this shortcoming, a simple extension of the proposed model is to replace the fixed $Z$ with a stochastic process $Z(\mathcal{T})$ with continuous sample trajectories. The stochastic process is chosen to be stationary with marginal distribution $Z(t)\sim p_{\text{noise}},\forall t\in\mathcal{T}$. This ensures that the generative model at any fixed $t$ is probabilistically equivalent with the fixed noise setting. Specifically, no changes to the training algorithm are needed since the training is based solely on time marginals. Allowing temporal stochasticity in the driving noise process results in a non-deterministic relationship between the states at different lead times, while we can keep the temporal consistency by ensuring that the driving process is sufficiently autocorrelated. A simple choice is to select $Z(\mathcal{T})$ as an Ornstein–Uhlenbeck (OU) process (Gillespie, 1996), $dZ(t)=-\rho Z(t)dt+\sqrt{2\rho}dW(t),$ for some correlation parameter $\rho>0$ and $W(t)$ being Brownian motion. This process can be easily simulated from to generate a noise sequence $\{Z(t_{i})\}_{i=1}^{N}$ for lead-times $\{t_{i}\}_{i=1}^{N}$, as done in Algorithm 2. The initial noise sample $Z(0)\sim\mathcal{N}(0,\mathbf{I})$ is perturbed according to the OU process by sampling new noise $\nu\sim\mathcal{N}(0,\mathbf{I})$ in each timestep. The update step in line 5 comes from the choice of an OU process for $Z(t)$ and corresponds to an exact simulation of the process at the specified lead times $\{t_{i}\}_{i=1}^{N}$, i.e., it guarantees the correct marginal distributions $z_{i}=Z(t_{i})\sim\mathcal{N}(0,\mathbf{I})$ for all lead times. An alternative to the OU process is to use a Gaussian process with stronger smoothness properties, e.g., with a squared exponential kernel. This could prove useful for formally proving time continuity of the resulting trajectories also under a stochastic noise process, but we leave such an analysis for future work.
@@
2:sample $\{z^{j}_{1}\}_{j=1}^{n_{\text{ens}}}\sim\mathcal{N}(0,\mathbf{I})$
@@
4:     sample $\{\nu^{j}_{i}\}_{j=1}^{n_{\text{ens}}}\sim\mathcal{N}(0,\mathbf{I})$
5:     $z_{i}^{j}\leftarrow\exp(-\rho(t_{i}-t_{i-1}))z^{j}_{i-1}+\sqrt{1-\exp(-2\rho(t_{i}-t_{i-1}))}\nu^{j}_{i}$
6:for all $i\in\{1,\dots,N\},j\in\{1,\dots,n_{\text{ens}}\}$ do $\triangleright$ Can be done fully in parallel for all $j$ and $i$
7:     $x^{j}_{i}\leftarrow\textsc{{Probability-Flow-Solver}}(z^{j}_{i},t_{i};x(\Omega),S_{\theta})$ return $\{x^{j}_{i}\}^{j=1:n_{\text{ens}}}_{i=1:N}$